In 1985, the Lincoln Memorial was 1 1/2 times as old as the Jefferson Memorial. If the Lincoln Memorial was built 21 years before the Jefferson Memorial, give the year each memorial was built.
Answer by scott8148(6628)

let x="year of Lincoln", so x+21="year of Jefferson" ... 1985-x=(3/2)(1985-(x+21))